x=5^1/4 5^-1/4 y=5^1/4 - 5^-1/4 prove that 5(x² y²)²=144?

1
Expert's answer
2022-02-07T13:55:13-0500

"x=5^{\frac14} +5^{-\frac14}, y=5^{\frac14} -5^{-\frac14}"

"x^2=5^{\frac12} +5^{-\frac12}+2\\
y^2=5^{\frac12} +5^{-\frac12}-2\\
x^2-y^2=4\\
(x^2-y^2)^2=16\\
5(x^2-y^2)^2=80"